a simple pendulum has a period of 2.4s on earth. the height of the bob is 0.5m. what will linear velocity of the bob be at the bottom of its swing

1
Expert's answer
2022-05-17T16:04:44-0400
"v=\frac{2\pi r}{T}=\frac{2\pi*0.5\:\rm m}{2.4\:\rm s}=1.3\:\rm m/s"